OpenStudy (anonymous):

If a donut shop sells a dozen donuts for $6.00, what would the cost of 9 donuts be?

hero (hero):

A dozen donuts is 12 donuts so

hero (hero):

\[\frac{12 \text{donuts}}{9 \text{donuts}} = \frac{$6}{$x}\] \[\frac{12}{9} = \frac{6}{x}\]

hero (hero):

Simplify and solve for x: \[\frac{4}{3} = \frac{6}{x}\] Cross multiply: 4x = 18 x = 18/4 x = 9/2 x = 4.5

hero (hero):

In terms of dollars, the cost of 9 donuts is $4.50

OpenStudy (unklerhaukus):

@Hero, please try to leave some of the work to the OP.

hero (hero):

1. The OP wasn't around 2. Someone posted an incorrect solution, so I wrote the correct solution. Then the other person deleted their solution and disappeared. Sorry to disappoint you.